Operations Report – April 13th

Crew 192 Operations Report 12Apr2018


Name of person filing report: Richard Blakeman

Non-nominal systems: 1. Refrigerator (resolved); 2. Kitchen sink faucet

Notes on non-nominal systems: The refrigerator appeared to have stopped working overnight. Outlet was inspected, fridge unplugged and plugged back in, has been working since. Will continue to monitor.

Sink is being worked on right now.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas. We have no knowledge of the gas or that 5 gallon container.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 925 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133428.9

Toilet tank emptied: No. Last emptied April 11, 2018.

ATV’s Used: Blue ATVs #1 and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 10%

# Hours the ATVs were Used today: 3

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, 2 one piece suits used. Radios working fine.

Summary of Hab operations: Cleaned and organized hab, including ceiling lights in the upper deck and various bins and shelves in lower deck. Organized food cabinets and cleaned additional kitchen appliances and pans. Fixed kitchen sink faucet.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank remains in the RAM for additional cleaning.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support:

– Regarding use of tables in the lower deck: There is a 3 foot wooden table in the spacesuit room and two 5 foot plastic tables in the lower deck. Requesting permission to swap the wooden table with one of the longer plastic ones so we can accommodate all one-piece EVA suits in the spacesuit room.

– Request permission to change out water filter that was delivered earlier today.

– Reviewed possible fixes for the skirt of the HAB as requested. Suggest using sheet metal screws and the electric drill to fasten skirting in place. Respectfully request permission to perform this repair via engineering EVA.

– There’s a picture attached of the base of a small tv screen/computer monitor that is in the hab, but does not seem to have the proper cords available for use. Does anyone know if the cords are available anywhere in the hab site?

– The laptop labeled “Hab” in the upper deck also does not have the proper cords to check for power. Please advise.

Thank you for removing the unneeded items we left outside. There are a few additional items we put outside for pickup.

Copyright © The Mars Society. All rights reserved. | Main Site